OpenStudy (anonymous):

Stumped on this one question: Any advice?: Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can?

OpenStudy (wolf1728):

I guess you'd first have to calculate the height of the can. volume = PI * r² * height 15 / PI * 1² = height height = 15 / PI = 4.77 Now you need the volume of a cone with height of 4.77 and radius of one. cone volume = (π • r² • h) / 3 cone volume = (3.14159 * 1 * 4.77) / 3 cone volume = 14.9853969576 / 3 cone volume = 4.995 or 5 rounded Here's a cone calculator to check your work. http://www.1728.org/volcone.htm

OpenStudy (anonymous):

Thank you! Couldn't get that one

OpenStudy (wolf1728):

Here is a quick solution I just noticed: cylinder volume = PI * r² * height cone volume = (PI * r² * height) / 3 The exact same numbers go in the exact same place so we can see that a cone of equal height and radius of a cylinder will have 3 times less volume due to the division by 3 in the cone formula.